The length of a rectanble is three inches more than the width. The area of the rectangle is 180 inches. Find the width of the rectangle.
I know that the length is x + 3
and the width is x
so the equation must be x(x + 3) = 180, but I don't know where I keep messing up when I go to solve it. Thanks so much for your help!!

Answer by Nate(3500)

You are on the right path:
x(x + 3) = 180
x^2 + 3x = 180
x^2 + 3x - 180 = 0
(x + 15)(x - 12) = 0
x = -15 and x = 12
Lenght can not be considered as negative ... x = 12
Length = 12 + 3 = 15 inches
Width = 12 inches